The Quasi-Polynomial Ansatz Applied to Counting Integer Partitions by Andrew V. Sills and Doron Zeilberger

References

Andrew V. Sills, DIMACS Center, 96 Frelinghuysen Road, Rutgers University, Piscataway, New Jersey 08854, USA and Georgia Southern University , asills [at] dimacs.rutgers.edu

Department of Mathematics, Hill Center- Busch Campus, Rutgers University, 110 Frelinghuysen Rd, Piscataway, NJ 08854-8019, USA, zeilberg at math dot rutgers dot edu